# Shock breakouts from compact CSM surrounding core-collapse SN progenitors may contribute significantly to the observed $\gtrsim 10$ TeV neutrino background

### v1

@@ 1 Introduction
Here, we discuss an alternative origin for the $\gtrsim 10$ TeV neutrino background – SN shocks breaking through compact optically thick CSM surrounding SN progenitors. Collisionless SN shocks driven into the winds surrounding massive stars and into the interstellar medium (ISM) have long been suggested as sources of CRs (Blandford & Eichler, 1987), and the high energy neutrino emission due to energy losses by these CRs was discussed by several authors (e.g. Murase et al., 2011; Petropoulou et al., 2017; Sarmah et al., 2022). The neutrino luminosity produced by shocks driven into typical wind/ISM extends over hundreds of days. It is too low to account for the neutrino background, and it is accompanied by high energy gamma-ray luminosity similar to the neutrino luminosity since the pair production optical depth is low for most of the emission, that occurs at large, $>10^{16}$ cm radii (see end of § 2.3 for a short discussion of the recent work of Sarmah et al., 2022; Murase, 2024).
We have shown (Katz et al., 2011, hereafter KSW11) that if a sufficiently optically thick CSM surrounds the SN progenitor star, then a large neutrino luminosity may be produced during and shortly after the SN shock breakout through the dense CSM shell, i.e. during and shortly after the shock transition from being radiation-mediated to collisionless (see also Li, 2019). Our motivation for considering the presence of such dense CSM was the observation of several X-ray flashes and low-luminosity gamma-ray bursts associated with SNe, which may be explained as (fast) shock breakouts from optically thick compact CSM. We return to this topic here since, over the past few years, the observational constraints on stellar mass loss preceding SN explosions have greatly improved, providing for the first time initial information on the prevalence and properties of mass loss in the type II SN progenitor population. While pre-explosion mass loss may also be expected in other rare types of core-collapse SN progenitors, we focus here on the population of type II progenitors, which are hydrogen-rich super-giants, for which observations allow one to draw initial constraints on the population as a whole.
@@ 2.1 Pre-explosion mass loss of type II SN progenitors
Independent evidence for the prevalent presence of optically thick CSM shells around spectroscopically regular type II SN progenitors is obtained from early, 1 day time scale observations of optical-UV SN light curves. Irani et al. (2024) carried out the first systematic analysis of early ($\sim 1$ d) optical-UV light curves of a large sample of type II SNe (see Morag et al., 2023, for a discussion of analyses of individual SNe). They find that while the early light curves of $\approx 50\%$ of type II SN are consistent with the emission from the expanding shocked stellar envelope, so-called “shock cooling” emission (see Waxman & Katz, 2017; Levinson & Nakar, 2020, for reviews), the light curves of the other 50% are inconsistent with shock cooling and indicate the presence of an optically thick CSM shell. The extended, days long, rise of the luminosity to high values of $10^{43}-10^{44}$ erg s-1 and the high color temperature, with most emission in the UV, are consistent with shock breakout from such a shell (Ofek et al., 2010; Waxman & Katz, 2017, and § 2.3): The shock breakout radius is inferred from the duration of the luminosity rise, which is given by the supernova driven shock crossing time, $R_{\rm br}=vt=10^{14.5}(t/3\,{\rm d})(v/10^{4}{\rm km{\rm\,s^{-1}}})$ where $v$ is the shock velocity; The shell mass within $R_{\rm br}$ is determined by the requirement that the optical depth at $R_{\rm br}$ equals $c/v$, $M=(c/v)4\pi R_{\rm br}^{2}/\kappa\approx 0.05M_{\odot}$; The observed breakout energy, $\approx 10^{49}$ erg is consistent with $0.5Mv^{2}=5\times 10^{49}$ erg; The high, $\sim 10$ eV color temperature is consistent with the breakout energy and radius. This work supports earlier suggestions from analyses of the rise time of Type II SN light curves (e.g. Förster et al., 2018; Morozova et al., 2018) that many of these explosions occur within a compact distribution of CSM.
@@ 2.3 Breakout neutrinos and photons
Our predicted neutrino luminosity is larger than predicted by Murase (2024) for wind interaction, largely because our model predicts larger shock energy deposited in the wind. For example, modeling SN 2003ixf with $\dot{M}/v_{\rm w,7}=0.01\,M_{\odot}$ yr-1 (their $D_{*}=0.1$) they obtain $\varepsilon_{\nu}^{2}dn_{\nu}/d\varepsilon_{\nu}=5\times 10^{45}$ erg, while Equation (9) predicts $7\times 10^{46}$ erg. This is due mainly to the low rate at which energy is deposited in the shell by the shock, $7\times 10^{42}\,{\rm erg\,s}^{-1}$ at $t=5$ d in their model, implying a bolometric luminosity $<3\times 10^{42}\,{\rm erg\,s}^{-1}$. In contrast, the observed luminosity is $\approx 3\times 10^{43}\,{\rm erg\,s}^{-1}$, consistent with the prediction of Equation (6). Future systematic observations of the early SN breakout shock emission will enable us to discriminate between models.
Finally, we note that Sarmah et al. (2022) predict a neutrino luminosity $\varepsilon_{\nu}^{2}d\dot{n}_{\nu}/d\varepsilon_{\nu}\approx 2\times 10^{40}\,{\rm erg\,s}^{-1}$ for $\approx 1$ yr, corresponding to $\varepsilon_{\nu}^{2}dn_{\nu}/d\varepsilon_{\nu}\approx 10^{48}\,{\rm erg}$, following type IIn supernovae. This extended emission is obtained assuming a $\dot{M}/v_{\rm w,7}\approx 0.1\,M_{\odot}$ yr-1 over 100 yr preceding the explosion, corresponding to $M_{\rm CSM}=10\,M_{\odot}$, large ejecta energy, $\gtrsim 10^{52}$ erg, and large bolometric luminosity $L\approx 10^{44}\,{\rm erg\,s}^{-1}$ lasting over a year. A possible example of such an event is SN 2010jl (e.g., Ofek et al., 2014b). However, such events are likely very rare [^1] Assuming one 2010jl like SN was found in 15 years to a distance of 50 Mpc over half the sky, then the rate of such events is in the range of $5\times 10^{-9}$ to $10^{-6}\,{\rm Mpc^{-3}yr^{-1}}$ (95% CL). , with a small rate compared to the overall type IIn rate, i.e. $\ll 10^{-5}\,{\rm Mpc^{-3}\,yr^{-1}}$, and cannot therefore contribute significantly to the neutrino background (see Equation (3)).
@@ 3 Discussion
A breakout producing $E_{\nu}(>1\,{\rm TeV})=10^{48}$ erg, corresponding to $M_{\rm CSM}=0.03M_{\odot}$ and electromagnetic emission of $2\times 10^{49}$  erg, will produce $>1$ muon induced neutrino events (on average) in a 1 km2 detector out to a distance of $\lesssim 5$ Mpc. Type II SNe are expected to occur within this volume at a rate of $\sim 1/20$ yr. For a 10 km2 detector, the rate of SNe producing $>1$ muon-induced neutrino events would increase to $\sim 1$/yr. An association by a 1 km2 detector of a single muon-induced 10 TeV neutrino with a very nearby SN ($\lesssim 10$ Mpc) within a few days of the explosion would be significant at approximately $99.9\%$ confidence level - for an atmospheric neutrino background rate of $\approx 2\times 10^{3}/$ yr (IceCube Collaboration et al., 2022) and arrival direction uncertainty of $<1$ squared degree, the fraction of the sky covered by the background neutrinos is $\approx 10^{-3}/(3\,{\rm d})$.
@@
E.O.O. is grateful for the support of grants from the Willner Family Leadership Institute, André Deloro Institute, Paul and Tina Gardner, The Norman E Alexander Family M Foundation ULTRASAT Data Center Fund, Israel Science Foundation, NSF-BSF, Israel Council for Higher Education (VATAT), Sagol Weizmann-MIT, and the Rosa and Emilio Segre Research Award. AGY’s research is supported by the André Deloro Institute for Space and Optics Research, the Center for Experimental Physics, a WIS-MIT Sagol grant, the Norman E Alexander Family M Foundation ULTRASAT Data Center Fund, and Yeda-Sela; AGY is the incumbent of the The Arlyn Imberman Professorial Chair.

### v3

@@ 1 Introduction
Here, we discuss an alternative origin for the $\gtrsim 10$ TeV neutrino background – SN shocks breaking through compact optically thick CSM surrounding SN progenitors. Collisionless SN shocks driven into the winds surrounding massive stars and into the interstellar medium (ISM) have long been suggested as sources of CRs (Blandford & Eichler, 1987), and the high energy neutrino emission due to energy losses by these CRs was discussed by several authors (e.g. Murase et al., 2011; Petropoulou et al., 2017; Sarmah et al., 2022). The neutrino luminosity produced by shocks driven into typical wind/ISM extends over hundreds of days. It is too low to account for the neutrino background, and it is accompanied by high energy gamma-ray luminosity similar to the neutrino luminosity since the pair production optical depth is low for most of the emission, that occurs at large, $>10^{16}$ cm radii (see the end of § 2.3 and § 3 for a short discussion of the recent work of Sarmah et al., 2022; Murase, 2024).
We have shown (Katz et al., 2011, hereafter KSW11) that if a sufficiently optically thick CSM surrounds the SN progenitor star, then a large neutrino luminosity may be produced during and shortly after the SN shock breakout through the dense CSM shell, i.e. during and shortly after the shock transition from being radiation-mediated to collisionless (see also Li, 2019). Our motivation for considering the presence of such dense CSM was the observation of several X-ray flashes and low-luminosity gamma-ray bursts associated with SNe, which may be explained as (fast) shock breakouts from optically thick compact CSM. We return to this topic here since, over the past few years, the observational constraints on stellar mass loss preceding SN explosions have greatly improved, providing for the first time initial information on the prevalence and properties of mass loss in the type II SN progenitor population. While pre-explosion mass loss may also be expected in other rare types of core-collapse SN progenitors, we focus here on the population of type II progenitors, which are hydrogen-rich super-giants, for which observations allow one to draw initial constraints on the population as a whole. The possible contribution of other SN types is briefly discussed in § 3.
@@ 2.1 Pre-explosion mass loss of type II SN progenitors
Independent evidence for the prevalent presence of optically thick CSM shells around spectroscopically regular type II SN progenitors is obtained from early, 1 day time scale observations of optical-UV SN light curves. Irani et al. (2024a) carried out the first systematic analysis of early ($\sim 1$ d) optical-UV light curves of a large sample of type II SNe (see Morag et al., 2023, for a discussion of analyses of individual SNe). They find that while the early light curves of $\approx 50\%$ of type II SN are consistent with the emission from the expanding shocked stellar envelope, so-called “shock cooling” emission (see Waxman & Katz, 2017; Levinson & Nakar, 2020, for reviews), the light curves of the other 50% are inconsistent with shock cooling and indicate the presence of an optically thick CSM shell. The extended, days long, rise of the luminosity to high values of $10^{43}-10^{44}$ erg s-1 and the high color temperature, with most emission in the UV, are consistent with shock breakout from such a shell (Ofek et al., 2010; Waxman & Katz, 2017, and § 2.3): The shock breakout radius is inferred from the duration of the luminosity rise, which is given by the supernova driven shock crossing time, $R_{\rm br}=vt=10^{14.5}(t/3\,{\rm d})(v/10^{4}{\rm km{\rm\,s^{-1}}})$ where $v$ is the shock velocity; The shell mass within $R_{\rm br}$ is determined by the requirement that the optical depth at $R_{\rm br}$ equals $c/v$, $M=(c/v)4\pi R_{\rm br}^{2}/\kappa\approx 0.05M_{\odot}$; The observed breakout energy, $\approx 10^{49}$ erg is consistent with $0.5Mv^{2}=5\times 10^{49}$ erg; The high, $\sim 10$ eV temperature is consistent with a blackbody radiation carrying the breakout energy at the breakout radius. This work supports earlier suggestions from analyses of the rise time of Type II SN light curves (e.g. Förster et al., 2018; Morozova et al., 2018) that many of these explosions occur within a compact distribution of CSM.
@@ 2.3 Breakout neutrinos and photons
We note that Sarmah et al. (2022) predict a neutrino luminosity $\varepsilon_{\nu}^{2}d\dot{n}_{\nu}/d\varepsilon_{\nu}\approx 2\times 10^{40}\,{\rm erg\,s}^{-1}$ for $\approx 1$ yr, corresponding to $\varepsilon_{\nu}^{2}dn_{\nu}/d\varepsilon_{\nu}\approx 10^{48}\,{\rm erg}$, following type IIn supernovae. This extended emission is obtained assuming a $\dot{M}/v_{\rm w,7}\approx 0.1\,M_{\odot}$ yr-1 over 100 yr preceding the explosion, corresponding to $M_{\rm CSM}=10\,M_{\odot}$, large ejecta energy, $\gtrsim 10^{52}$ erg, and large bolometric luminosity $L\approx 10^{44}\,{\rm erg\,s}^{-1}$ lasting over a year. A possible example of such an event is SN 2010jl (e.g., Ofek et al., 2014b). However, such events are likely very rare [^1] Assuming one 2010jl like SN was found in 15 years to a distance of 50 Mpc over half the sky, then the rate of such events is in the range of $5\times 10^{-9}$ to $10^{-6}\,{\rm Mpc^{-3}yr^{-1}}$ (95% CL). , with a small rate compared to the overall type IIn rate, i.e. $\ll 10^{-5}\,{\rm Mpc^{-3}\,yr^{-1}}$, and cannot, therefore, contribute significantly to the neutrino background (see Equation (3)).
@@ 3 Discussion
A breakout producing $E_{\nu}(>1\,{\rm TeV})=10^{48}$ erg, corresponding to $M_{\rm CSM}=0.03M_{\odot}$ and electromagnetic emission of $2\times 10^{49}$ erg, will produce $>1$ muon induced neutrino events (on average) in a 1 km2 detector out to a distance of $\lesssim 5$ Mpc. Type II SNe are expected to occur within this volume at a rate of $\sim 1/20$ yr. For a 10 km2 detector, the rate of SNe producing $>1$ muon-induced neutrino events would increase to $\sim 1$/yr. An association by a 1 km2 detector of a single muon-induced 10 TeV neutrino with a very nearby SN ($\lesssim 10$ Mpc) within a few days of the explosion would be significant at approximately $99.9\%$ confidence level - for an atmospheric neutrino background rate of $\approx 2\times 10^{3}/$ yr (IceCube Collaboration et al., 2022) and arrival direction uncertainty of $<1$ squared degree, the fraction of the sky covered by the background neutrinos is $\approx 10^{-3}/(3\,{\rm d})$. The current upper limits on the energy emitted in high energy neutrinos following SNe ($3\times 10^{48}$ erg and $10^{49}$ erg for SNe of type IIP and IIn respectively, Abbasi et al., 2023), are larger than, but close to, the emission predicted in our analysis (see Eq. (10)). These limits are, however, not directly applicable to the scenario that we consider: They were obtained for neutrino emission over long time windows (100 d and longer) starting at the time of first optical SN detection, which is typically later than the few days time scale over which most of the neutrino emission is predicted to occur.
@@
The neutrino luminosity predicted in Murase (2024) for SN 2003ixf is significantly lower than predicted here mainly because they adopt CSM density and shock velocity that yield a substantially lower rate of shock energy deposition in the CSM, $7\times 10^{42}\,{\rm erg\,s}^{-1}$ compared to our $\approx 10^{44}\,{\rm erg\,s}^{-1}$ (see Eq. (6); the discrepancy in total fluence is smaller). Our parameter choice is based on observations, including in particular the observed $\approx 3\times 10^{43}\,{\rm erg\,s}^{-1}$ luminosity, that were not available at the time of their paper submission. When similar values are adopted for the CSM density and shock velocity (and CR acceleration efficiency), our neutrino luminosity estimate is similar to what would be obtained by Murase (2024).
While our discussion has focused on type II SNe, which is the most common type of events, recent evidence shows that other types of massive star explosions also occur within compact CSM, including some type Ic events lacking hydrogen and helium (Irani et al., 2024b), as well as rare populations of events where the CSM is dominated by helium (Type Ibn, Pastorello et al., 2007, 2008; Karamehmetoglu et al., 2021; Hosseinzadeh et al., 2017), or carbon and oxygen (type Icn, Gal-Yam et al., 2022; Perley et al., 2022; Pellegrino et al., 2022). The equations derived in § 2.3 for the neutrino emission are valid for all progenitors surrounded by hydrogen-dominated CSM and are straightforward to generalize to He- and C/O-dominated CSM. While the rate of type Ibc SNe is $\approx 1/3$ the type II SN rate, their contribution to the detected cosmic high-energy neutrino signal may be significant if compact dense CSM is common for such explosions, noting particularly that they may be characterized by higher breakout velocities due to the more compact progenitor structure. Further observations are required to constrain the CSM properties of this population.
@@
We thank K. Murase for useful discussions and the anonymous referee for significant constructive comments. E.W.’s research is partially supported by grants from the André Deloro Institute for Space and Optics Research, the Norman E Alexander Family M Foundation ULTRASAT Data Center Fund, and the Sagol Weizmann-MIT Bridge Program. E.O.O. is grateful for the support of grants from the Norman E Alexander Family M Foundation ULTRASAT Data Center Fund, and the Willner Family Leadership Institute for the Weizmann Institute of Science, Willner Family Leadership Institute, André Deloro Institute, Paul and Tina Gardner, Israel Science Foundation, NSF-BSF, Israel Council for Higher Education (VATAT), Sagol Weizmann-MIT, and the Rosa and Emilio Segre Research Award. AGY’s research is supported by the André Deloro Institute for Space and Optics Research, the Center for Experimental Physics, a WIS-MIT Sagol grant, the Norman E Alexander Family M Foundation ULTRASAT Data Center Fund, and Yeda-Sela; AGY is the incumbent of the The Arlyn Imberman Professorial Chair.